Iron exists as iron ore in Earth’s crust as iron(III) oxide. To obtain pure iron, the ore must be “smelted” using a chemical reaction. Modern iron smelting combines carbon monoxide with the iron ore at temperatures exceeding 1000 ०C in a series of reactions that results in (unbalanced):

Fe2O3 (s) + CO (g) ➝ Fe (l) + CO2 (g)


Josiah, a chemical engineer, is calibrating the equipment used for smelting at the Iron Ore Company of Canada in Labrador City, Newfoundland and Labrador. Because the smelting process produces carbon dioxide as a byproduct, Canadian climate law dictates that he must measure the amount of carbon dioxide produced by his plant. The standards are reported in terms of molecules of CO2 output per day. If Josiah's plant produces an average of 1.500 x 1028 molecules of CO2 per day, calculate the mass of iron ore smelted by the plant per day.
1
Expert's answer
2020-05-26T14:21:43-0400

Josiah's plant produces an average of 1.500 x 1028 molecules of CO2 per day.

n(CO2) = N/Na = 1.500 x 1028 molecules / 6.02 x 1023 mol-1 = 0.2492 x 105 mol

Fe2O3 (s) + 3CO (g) ➝ 2Fe (l) + 3CO2 (g) 

According to the equation:

n(Fe) = 0.2492*105*2/3 = 0.1661 x 105 mol

The mass of iron ore smelted by the plant per day:

m(Fe) = n*M = 0.1661 x 105 mol*56g/mol = 9.3 x 105 g = 9.3 x 102 kg